### v1

# Evaluation Should Not Ignore Variation: On the Impact of Reference Set Choice on Summarization Metrics

Abstract
Human language production exhibits remarkable richness and variation, reflecting diverse communication styles and intents. However, this variation is often overlooked in summarization evaluation. While having multiple reference summaries is known to improve correlation with human judgments, the impact of using different reference sets on reference-based metrics has not been systematically investigated. This work examines the sensitivity of widely used reference-based metrics in relation to the choice of reference sets, analyzing three diverse multi-reference summarization datasets: SummEval, GUMSum, and DUC2004. We demonstrate that many popular metrics exhibit significant instability. This instability is particularly concerning for n-gram-based metrics like ROUGE, where model rankings vary depending on the reference sets, undermining the reliability of model comparisons. We also collect human judgments on LLM outputs for genre-diverse data and examine their correlation with metrics to supplement existing findings beyond newswire summaries, finding weak-to-no correlation. Taken together, we recommend incorporating reference set variation into summarization evaluation to enhance consistency alongside correlation with human judgments, especially when evaluating LLMs.
Evaluation Should Not Ignore Variation:
On the Impact of Reference Set Choice on Summarization Metrics
[…]
1 Introduction
Human-written texts vary widely in terms of length, style, communicative intent, lexical/syntactical choices, and numerous other dimensions Giulianelli et al. (2023); Liu and Zeldes (2023); Rezapour et al. (2022); Baan et al. (2023). Such variation poses a significant challenge in the evaluation of summarization systems Lloret et al. (2018); Celikyilmaz et al. (2021). Traditional summarization metrics typically rely on comparing system outputs to one or more references, treating these references as a “gold standard”. Although the limitations of reference-based metrics have long been acknowledged Rankel et al. (2013); Louis and Nenkova (2013); Reiter (2018); Peyrard (2019); Fabbri et al. (2021); Goyal et al. (2023),
Figure 1: Human-written summaries are diverse. Using a human-written reference instead of another makes evaluation metrics fluctuate, and affects model ranking.
they remain widely popular due to their simplicity, low compute requirements, relative ease of adaptation to different languages, and reproducibility. In 2024, out of $21$ ACL papers mentioning “summarization” in their title, $19$ ($90$%) include a reference-based metric in their evaluation, with ROUGE Lin (2004) being the most common ($71$%), followed by BERTScore (Zhang et al. 2020, $52$%).
The assumption behind the use of reference-based metrics is that systems outputs that are more similar to the reference(s) are better, due to their “human-likeness” Gehrmann et al. (2023). However, the significant variation in human-written summaries implies that evaluating system outputs against a single or limited set of references has inherent drawback. Previous research has extensively looked at correlation between metrics and human judgments in summarization, further exploring the use of multiple references to improve such correlation Lin (2004); Belz and Reiter (2006); Fabbri et al. (2021); Tang et al. (2024). However, a much less studied question is the extent to which automatic metrics are sensitive to the choice of human-written reference summaries, as shown in Figure 1. In other words, are these metrics stable across different plausible gold-standard references? If metric scores vary significantly with the selected reference(s), this variation calls into question the reliability of many evaluation practices in the field.
In this work, we quantify the impact of reference choice on automatic evaluation metrics for summarization. Our contributions are as follows:
  1. [1]
    We investigate how different reference sets affect system rankings. Our results show that system rankings based on n-gram-matching metrics (e.g., ROUGE) strongly depend on the choice of the reference(s), undermining the reliability of model comparisons. However, rankings based on more semantically-oriented metrics exhibit greater stability.
  2. [2]
    We examine the robustness of widely-used reference-based metrics at the instance and dataset level. Our analysis reveals that the variation in scores introduced by the choice of reference on a dataset often exceeds the variation observed in state-of-the-art (SOTA) summarization models.
  3. [3]
    We collect new human judgment scores on Large Language Model (LLM) outputs for the genre-diverse GUMSum Liu and Zeldes (2023) dataset. We use these data to reassess the correlation between automatic metrics and human judgments, complementing earlier SummEval evaluation (Fabbri et al., 2021) limited to pre-LLM models and newswire data. We find that correlations tend to increase with the number of references, and that the metric with the highest correlation varies depending on the evaluation dimension and the number of references.
Our analysis reveals that few metrics tend to show reasonable correlation with human judgments and robustness to the reference sets, especially when scoring LLM outputs.
2 Related Work
Summarization Evaluation.
Recent advances in Natural Language Generation (NLG) have significantly enhanced the development of automatic summarization systems. However, their evaluation remains an open problem Celikyilmaz et al. (2021); Goyal et al. (2023). Summarization evaluation metrics are broadly categorized into reference-based and reference-free Lloret et al. (2018). Reference-based metrics compare system outputs to human-written reference summaries, relying on methods such as n-gram overlap Lin (2004); Papineni et al. (2002), embedding similarity Ng and Abrecht (2015); Zhao et al. (2019); Zhang et al. (2020), or model-based evaluation techniques Peyrard et al. (2017); Scialom et al. (2019); Yuan et al. (2021). In contrast, reference-free summarization metrics do not assume a gold standard Yuan et al. (2021); Vasilyev et al. (2020); Gao et al. (2020). Lastly, growing research starts to leverage LLMs as evaluators, with or without references Liu et al. (2023); Song et al. (2024); Li et al. (2024).
Metrics Meta-Evaluation.
Meta-evaluation of summarization metrics typically focuses on the extent to which they can be used as a proxy for human evaluation. Reiter and Belz (2009) examined the validity of automatic scores for NLG tasks, while Rankel et al. (2013) focused on ROUGE and its correlation with human judgments. Peyrard (2019) showed that metrics with reasonable correlation on lower-quality outputs tend to diverge when output quality increases. Caglayan et al. (2020) demonstrated various idiosyncrasies of automatic evaluation metrics, noting that high correlation with human judgments is not sufficient to characterize their reliability. Fabbri et al. (2021) performed a large-scale meta-evaluation of summarization metrics, and found that most metrics have low correlation with human judgments on coherence, while relevance is weakly or moderately correlated.
While most existing research focused on correlation to human scores, Tang et al. (2024) addressed the challenge of evaluation when a limited number of references is available. They proposed leveraging LLMs to diversify the references, expanding the evaluation coverage and improving the correlation with humans. Their results show that increasing the number of references significantly enhances the reliability of existing evaluation metrics in terms of correlation. However, since LLM outputs tend to show less variability and follow distinct patterns compared to human-produced content Giulianelli et al. (2023); Guo et al. (2024); Shur-Ofry et al. (2024); Reinhart et al. (2025), relying on them to replace human references might introduce biases.
dataset #doc genre references outputs
#sums #chars #toks model outputs human judgments
SummEval 100 news 1+10 226.3 43.1 Fabbri et al. (2021) + 4 LLMs Fabbri et al. (2021)
GUMSum 48 12 genres 5 291.3 52.1 4 LLMs collected in this work
DUC2004 489 news 4 70.0 11.9 4 LLMs n/a
Table 1: Multi-reference summarization datasets. #sums indicates the number of human-written references per instance. We generate outputs using four LLMs and collect a new set of human judgments for GUMSum.
3 Experimental Setup
To quantify the impact of different human-written references on the scores of automatic metrics, we exam multiple elements. For datasets, we use SummEval Fabbri et al. (2021), GUMSum Liu and Zeldes (2023), and DUC2004 Dang and Croft (2004), which contain multiple human-written summaries (§3.1), to assess how different reference summaries affect metric performances. Next, to assess summarization models, we use the existing outputs provided by Fabbri et al. (2021) for SummEval. As these outputs predate LLMs, we additionally collect outputs using LLMs (§3.2) for all three datasets. Lastly, to compute the correlations with humans, we use the human judgments available in SummEval and gather new human ratings for GUMSum on both human and LLM-generated summaries (§3.3). We prioritize GUMSum over DUC2004, as it includes multiple genres beyond news data. We comply with the license of the existing datasets. For newly collected model outputs and human judgments, we follow the license of the corresponding underlying datasets. Our metric selection is outlined in §3.4.
3.1 Human-written Summaries
Table 1 provides an overview of the three datasets.
SummEval Fabbri et al. (2021) is built on top of CNN/DM Hermann et al. (2015); Nallapati et al. (2016), containing news articles and human-written highlights. Authors selected $100$ instances from the test set; for these, in addition to the highlight-based summary in CNN/DM, ten references were crowd-sourced Kryscinski et al. (2019).
GUMSum Liu and Zeldes (2023) contains summaries created following general and genre-specific guidelines [^1] https://wiki.gucorpling.org/gum/summarization to function as a substitute for the source Nenkova and McKeown (2011). We focus on the $48$ documents in the dev and test sets, which contain five human-written summaries each Lin and Zeldes (2025), evenly distributed across $12$ genres.
DUC2004 Task1 Dang and Croft (2004) consists of $489$ news documents, most with four references. The guidelines allow the summaries to be in the form of short sentences or lists of keywords. [^2] https://duc.nist.gov/duc2004/tasks.html DUC2004 references are thus extremely concise (only up to $75$ characters). The dataset has played a significant role in summarization research and was part of the annual TREC conference evaluation.
3.2 Model Outputs
Fabbri et al. (2021) collected model outputs for SummEval from $24$ extractive and abstractive summarization systems, which were SOTA between 2017 and 2019. We focus on the $16$ models for which they provided human judgments.
For all datasets, we also include summaries generated by contemporary LLMs. This is crucial given that prior studies have demonstrated that evaluation metrics often show lower correlation with high-quality outputs Peyrard (2019); Alva-Manchego et al. (2021). We see a similar pattern for LLMs (§4.4). For consistency purposes, we follow Lin and Zeldes (2025) and use Llama3-3B-Instruct Hermann et al. (2015), Qwen-2.5-7B-Instruct Qwen et al. (2025), Claude-3.5 Anthropic (2024), and GPT-4o OpenAI (2024). For each LLM, we generate a single summary. This way, we emphasize LLM variety over multiple generations. Details on the generation parameters and prompts etc. are reported in Appendix A.
3.3 Human Judgments
SummEval (Fabbri et al., 2021) contains expert judgments that assess summaries based on four criteria: coherence, consistency, fluency, and relevance, using a Likert scale of $1$-$5$ Likert (1932).
To measure how well automatic metrics align with human judgments in different genres, and to study whether findings on pre-LLM models align with those on LLM outputs, we conduct a human evaluation on the $48$ GUMSum documents. We recruited three Master’s students in Computational Linguistics and instructed them to evaluate four LLM outputs (§3.2) and five human references, following Fabbri et al. (2021)’s criteria.
Summarizer Coh. Con. Flu. Rel. best worst
claude 4.750.45 4.480.65 4.820.42 4.260.71 0.170.38 0.030.18
gpt4o 4.420.61 4.610.63 4.740.45 4.070.65 0.110.32 0.120.32
Qwen 4.660.56 4.330.84 4.680.52 4.230.75 0.160.37 0.120.33
Llama3 4.710.51 4.140.97 4.780.42 4.090.86 0.120.32 0.200.40
humans 4.540.57 4.480.71 4.700.51 4.220.69 0.090.28 0.100.31
Table 2: Human judgments on GUMSum for system-generated versus human-written summaries. best and worst indicate the percentage of evaluators who voted the summarizer as best/worst averaged over documents.
LLM-generated and human-written summaries were anonymized and shuffled. We also asked the evaluators to pick one best and one worst summary for each document. Table 2 reports the results. We note that overall, Claude scored the best. GPT-4o gets the highest consistency but the lowest coherence and relevance, and thus the least picked LLM output. Interestingly, LLM outputs typically receive higher scores than human-written references.
3.4 Evaluation Metrics
We examine the following reference-based metrics chosen due to their widespread use. All metric ranges fall in 00-$100$. Appendix B provides details.
ROUGE Lin (2004) is the most popular metric for summarization. ROUGE-N computes n-gram overlap between a hypothesis and the references. ROUGE-L leverages the longest common subsequence, accounting for the word order. When evaluating with multiple references, ROUGE considers either the maximum or the mean of the n-gram overlap (ROUGEmax and ROUGEavg). We report the F1-score.
BLEU Papineni et al. (2002) is an n-gram overlap metric primarily used to assess translations. It is precision-based and incorporates a brevity penalty. When multiple references are provided, the n-gram count is clipped at the maximum count of n-grams in a single reference and the length of the reference closest in size to the hypothesis is considered.
METEOR Banerjee and Lavie (2005) incorporates multiple linguistic aspects, including synonym matching, stemming, and word order, making it more robust in capturing semantic equivalence. While primarily designed for translation, it has also been used to assess summaries. With multiple references, the maximum score is considered.
BERTScore Zhang et al. (2020) leverages pre-trained contextual embeddings and considers the cosine similarity between the embeddings of the hypothesis and the reference tokens. With multiple references, the final score is the maximum among the individual scores. We report the F1 score.
BLEURT (Sellam et al., 2020) is a model-based metric that leverages a pre-trained BERT fine-tuned on human judgments. The metric is not designed to handle multiple references; thus, we compute individual scores for each reference and consider the maximum value.
4 Reference Variability and Metrics Robustness
Figure 2: Variation in human-written summaries across datasets, measures inspired by Giulianelli et al. (2023).
Reference-based metrics assume that more human-like outputs deserve higher scores. However, human summaries are very diverse. This section examines how metrics fluctuate with different human references. By analyzing metric robustness, we aim to understand how conclusions about models, drawn from reference-based metrics, might change when different sets of human-written references are used, thereby undermining evaluation reliability.
4.1 Human-written Summaries are Diverse
Human-written summaries show substantial diversity. We assess the variability in the multi-reference datasets following Giulianelli et al. (2023). For each pair of human-written summaries for the same instance in the three datasets, we report the lexical similarity (the overlapping distinct n-grams between two strings), the syntactic similarity (the overlap of part-of-speech tag n-grams), and the semantic similarity (the cosine and euclidean similarity between the embeddings of the two strings).
Figure 2 shows these variations. At the dataset level, DUC and SummEval show the lowest similarity among human-written summaries across all dimensions. For GUMSum, summaries are more similar to each other. We hypothesize that this is likely due to the constrained annotation guidelines. [^3] https://wiki.gucorpling.org/gum/summarization It is also worth noting that the similarities revealed here are between different human-written summaries for a given instance as opposed to summaries across genres, for which we still expect significant variations, as demonstrated by Liu and Zeldes (2023). Overall, summaries tend to be similar at the syntactic level, less so at the semantic and lexical level. We also observe that LLM outputs show lower diversity (Appendix C), consistently with previous work Giulianelli et al. (2023).
4.2 Automatic Metrics Fluctuate Substantially at the Instance Level
Given the diversity in human-written summaries, we aim to quantify metric fluctuations at the instance level when using a different set of human-written references. For an automatic evaluation metric $M$ and a set of human-written references $R=\{r_{1},r_{2},\dots,r_{N}\}$, we compute $M(r_{i},R-\{r_{i}\})$. In other words, for each document, we score each human-written summary using all the others as the reference set. Figure 3 exemplifies the observed variability at the instance level measured by ROUGE-Lavg on the three datasets. For SummEval, we also highlight the scraped original reference in the CNN/DM dataset with a cross. The quality of these scraped references versus the ten crowd-sourced ones is discussed further in Appendix D.
Figure 3: Instance-level variation for ROUGE-Lavg. For every document (shown first $30$, one per line), we plot the score for every human-written reference against all other references (using the same color per source to aid interpretation). The original CNN/DM reference in SummEval is marked by a cross.
Scores assigned to human-written summaries are often low.
For example, the averaged ROUGE-Lavg scores are $28.52_{\pm 5}$, $27.46_{\pm 3}$, $24.88_{\pm 5.3}$ for SummEval, GUMSum, and DUC2004. Given the assumption that human reference summaries are of high quality (i.e., “gold”), metrics should produce high scores. Instead, they do not typically reflect this property.
Figure 4: Ranges of variability at the instance level on SummEval. For each instance, we compute the range of the scores of the references against the remaining ones. The trends for ROUGEmax and ROUGEavg are similar.
Human-written references scores vary widely.
Figure 4 summarizes the instance-level variability of the individual scores (in Figure 3) for all evaluation metrics on SummEval (corresponding figures for GUMSum and DUC are in Appendix E). For each metric, we compute the range (i.e., the difference between the maximum and the minimum score) when scoring human-written references against all the others ($M(r_{i},R-\{r_{i}\})$). Figure 4 shows the histogram of such ranges. Note that the ranges of variation observed within human-written references are on average very high.
Understanding the magnitude of such a range might not be obvious. For instance, an increase of $10$ points of BERTScore (which typically returns scores compressed in the high range of the scale) might indicate a much larger improvement in performance than an increase of $10$ points of ROUGE-1.444The relative dynamics between metrics have been studied by Kocmi et al. (2024) in machine translation. To contextualize the magnitude of variation for each metric, we also report the range of performance of summarization systems. Thus, for a model $S$, given its output $o_{i}$ for instance $i$, we score it through $M(o_{i},R)$. Although these values are not directly comparable and should be interpreted with caution due to the use of different reference sets, they help contextualize the magnitude of the results and its potential impact on evaluation. For example, ROUGE-1max assigned to human-written references vary by about $35$ points on average (the green dashed line in Figure 4). In contrast, the mean range is less than $20$ points across all model outputs (orange line), and much lower for LLM-generated outputs (blue line). These findings highlight the significance of the observed variability and suggest that the ranking of summarization models is highly sensitive to the reference set.
4.3 System Ranking Depends on the Reference(s) for n-gram-based Metrics
Figure 5: Rank stability when increasing the number of references. ROUGEmax is presented. Note that we use different ranges for the y axis for each dataset to improve readability.
While we observed variability at the instance level, summarization metrics are typically designed to evaluate models across entire datasets, rather than individual instances. In this section, we investigate to what degree standard summarization metrics can handle the variability observed in human-written references when ranking summarization systems.
Procedure.
We sample $k$ humans-written references ($k\in[1,N]$, where $N$ is the number of references for each document) from all available references for each instance. We then score the outputs of each summarization system using the same set of references. Given $M$ systems $S_{a}$, $S_{b}$, …, $S_{M}$, the metric induces a ranking $S_{a}\succ S_{b}\succ\dots\succ S_{M}$. This process is repeated $100$ times, yielding $100$ resulting rankings. We compute the pairwise Kendall rank correlation coefficient Kendall (1938) between such ranks. High correlation indicates that different sets of references lead to similar model ordering, even when different sets of references are used. Figure 5 reports the average correlation for pairs of ranks for each dataset and metric, from using a $k$ human-written summaries as references. ROUGEmax is shown in Figure 5, and ROUGEavg is reported in Figure 11 in Appendix F.
Single Reference.
Evaluating with a single reference is common in summarization, as most datasets provide only one human-written summary. Figure 5 (looking at $k=1$) shows the stability of different metrics with a single reference across the three datasets. We find that BLEU and ROUGE have very weak to moderate correlation between ranks across different references. In other words, using two different sets of plausible references would likely lead to different conclusions on relative model performance. We also notice a large variability among the individual pairs of rankings, with some showing negative correlation (refer to Table 4 in Appendix F for results on individual metrics and datasets).
In contrast, more semantically-oriented metrics show a greater stability. For SummEval, BLEURT shows the highest correlation between ranks, followed by METEOR and BERTScore. BLEURT and METEOR confirm their stability on GUMSum when ranking the LLM outputs. Other metrics (including BERTScore) show low or no correlation on GUMSum, with the exception of ROUGE-1. In all cases, metrics show much higher stability on DUC, for which all average correlations are above $0.7$. We speculate that the high stability might be due to an artifact introduced by the short length of the summaries, required by the guidelines.
In summary, n-gram-matching metrics, though simple, are highly reference-dependent, undermining consistent model evaluation, while semantically-oriented ones show greater stability. Therefore, we recommend always using model-based metrics in benchmarks with a single reference. When cost is a factor, METEOR might offer a good balance of stability and affordability.
Multiple References.
When scoring model outputs against a set of $k>1$ randomly sampled references, we observe that the correlation between rankings obtained with different human-written references generally improves when increasing the number of references. This increased stability is expected and in line with similar findings that associate a larger number of references with a higher correlation with humans Lin (2004).
However, the stability varies by metric. ROUGE (especially ROUGEmax) and BLEU tend to have lower correlation between ranks than other metrics. As an example, the ROUGEmax scores require $5$-$10$ references to reach a level of stability that is comparable to that of BERTScore on SummEval with a single reference. ROUGEavg has a better stability than ROUGEmax, especially with a larger set of references. For example, on SummEval, ROUGE-Lavg has higher stability than BERTScore for $k>3$, while on GUMSum, ROUGE-2avg is the second most stable metric for $k>3$. On all datasets, BLEURT and METEOR remain stable even with a single reference, with METEOR showing remarkable stability despite its simplicity.
In general, trends on SummEval are clearer and simpler to interpret than the other two datasets. We speculate that this is due to the larger number of models used ($16$ pre-LLM models+$4$ LLMs on SummEval vs $4$ LLMs on GUMSum and DUC). BLEURT, METEOR, and BERTScore show the highest stability, while n-gram-based metrics show low to average correlation between ranks even when multiple references are used. The cases of GUMSum and DUC2004 are more complex to interpret and might be less meaningful given fewer model outputs (i.e., only four LLM outputs). For GUMSum, BLEURT continues to show high inter-rank correlation, with METEOR being the second most stable. BERTScore, on the other hand, shows poor stability. Similar to the case with $k=1$, on DUC2004, all metrics show high stability, likely due to summaries being very short, as dictated by the guidelines.
4.4 Correlation with Human Judgments
Figure 6: Pearson correlation at the instance level on SummEval (top) and GUMSum (bottom).
In addition to stability, automatic metrics should correlate with human judgments. We compute correlations for SummEval and GUMSum, for which we have human judgments, [^5] For SummEval, we use the $16$ models studied by Fabbri et al. (2021); for GUMSum, the four LLMs. at the instance and system level as the number of references $k$ increases. [^6] For example, when considering two references, we consider the sets ($\binom{N}{2}$) of human-written references, where $N$ is the total number of references. We compute the scores using such references as gold standard, and report the mean.
Instance-level Correlation.
Figure 6 reports the instance-level correlation for SummEval (top) and GUMSum (bottom), respectively, versus number of references. We show ROUGEmax; corresponding figures using ROUGEavg are in Appendix G.
We notice weak-to-no correlation on both datasets. All correlations are generally higher on SummEval (where we consider outputs from the pre-LLM era) than on GUMSum (where we consider LLMs), in accordance to previous work showing that correlation with human judgments decreases as the quality of the outputs improves Peyrard (2019). For SummEval, increasing the number of references consistently leads to better correlation. This effect vanishes on GUMSum, where a larger reference set leads to no effect or slightly lower correlation. For SummEval, BERTScore shows the highest correlation on all dimensions but consistency, for which METEOR and ROUGEavg are a better proxy. Notice how the best metric in terms of correlation with human judgment depends on the considered criterion and the available number of references: BLEURT, for example, typically has low correlation when considering one reference only, performing worse than ROUGE. However, its performance improves when more references are considered, surpassing the scores of n-gram-based metrics.
System-level Correlation.
System-level correlation is generally higher than instance-level correlation on SummEval; however, many criteria still show weak to moderate correlation when one or very few references are included. In most cases, such correlation tends to improve with the number of references. This is not the case for ROUGEmax, especially when considering consistency. The full results are provided in Figure 13 in the Appendix G. GUMSum is excluded from this analysis due to the small number of systems available.
5 Conclusions
In this work, we investigate how reference sets impacts the reliability of reference-based summarization metrics. Our analysis across three multi-reference datasets—SummEval, GUMSum, and DUC2004—reveals that token-matching metrics such as ROUGE and BLEU are highly sensitive to the reference(s). This sensitivity leads to instability in system rankings, particularly when only a small number of references are available. We thus recommend avoiding such metrics, complementing previous work Schmidtova et al. (2024).
We find that increasing the number of reference summaries consistently improves both the stability of metric scores and their alignment with human judgments. In these conditions, n-gram-based metrics such as ROUGE-N become more reliable. Moreover, ROUGEavg should typically be preferred over ROUGEmax. However, when few references are available—especially in domains with high output variability—model-based metrics should be preferred, as they offer greater robustness despite their higher computational costs.
Our findings highlight the need to incorporate reference set variation into evaluation frameworks and suggest that future metric development should explicitly account for this dimension. We thus advocate for the creation of larger, more diverse multi-reference datasets, as well as for metric designs that are inherently robust to reference variability. Such efforts will be key to ensuring fairer and more reliable and human-aligned evaluation practices in summarization in the LLM era and beyond.
Limitations
While our study highlights challenges posed by reference set variation in summarization evaluation, it also comes with several limitations.
Firstly, although we focus on multi-reference datasets—SummEval, DUC2004, and GUMSum—such datasets remain relatively small. This reflects a broader limitation in current meta-evaluation practices, where multi-reference resources are the norm despite their limited scale.
Secondly, our analysis primarily targets standard evaluation criteria such as coherence, consistency, fluency, and relevance. While these are widely adopted and established, they do not capture all the nuances of summary quality, especially when the source texts are genre-diverse. More fine-grained human annotations and task-specific dimensions (e.g., factuality for news, stance for opinion pieces) would allow a deeper understanding of metric behavior under reference variation.
Thirdly, while we assess metric robustness using four systems (all of which are LLMs as opposed to the case for SummEval, where $16$ pre-LLM supervised systems are available) on the GUMSum and DUC2004 datasets, the relatively small number of models limits the generalization of the conclusions we can draw about system-level stability. Future work should include a larger and more diverse set of systems to better assess generalizability.
Lastly, despite our efforts to use contemporary LLMs, there remains a potential risk of data contamination. Since some of these datasets may have been seen during pretraining, this could affect both the outputs generated by LLMs and their evaluation scores. While we do not observe signs of memorization, we acknowledge that further controlled experiments are necessary to rigorously assess this risk.
Ethics Statement
All annotators involved in the collection of human judgments for the GUMSum dataset have given their consent to participate in the study and to allow us to publish the collected scores. They were paid according to national standards.
We used AI-assistants for improving the clarity of the text and of the figures.
Acknowledgments
We would like to thank the members of the MaiNLP lab for their valuable feedback, especially to Andreas Säuberli, Diego Frassinelli, Shijia Zhou, Soh-Eun (Ryan) Shim, and Domenico De Cristofaro. We also thank Pingjun Hong, Yuchen Mao, and Sebastian Loftus for their annotations. We recognize the support for Yang Janet Liu and Barbara Plank through the ERC Consolidator Grant DIALECT 101043235.
Appendix A LLM-generated Summaries
For each source, we generate a single summary using each of the four LLMs. For SummEval, the sources corresponding to the $100$ multi-reference summaries are taken into account. For DUC2004 Task 1, we generate summaries for the whole dataset ($489$ instances). For GUMSum, we focus on the dev and test set (in total $48$ instances). Thus, we generate a total of $2548$ summaries.
A.1 Prompts
A.1.1 SummEval
Article: {full_text}. Summarize the article in three sentences. Summary:
A.1.2 DUC2004 Task 1
The task is to create a very short single-document summary for the article below.
A very short summary should not be longer than 75 characters - this includes spaces and punctuation.
We will chop off characters beyond the 75th, so please do not include more than 75.
A very short summary could look like a newspaper headline, be a list of important terms or phrases separated by commas, a sentence, etc.
It should not contain any formatting, i.e., no indented lists, etc. Feel free to use your own words.
Article: {full_text} Summary:
A.1.3 GUMSum
Following Liu and Zeldes (2023), a general prompt was used to instruct LLMs to generate summaries, as shown below. Summarize the following article in 1 sentence. Make sure your summary is one sentence long and does not exceed 380 characters. Example of summary style: example
{doc_text}
Summary:
A.2 LLM Output Evaluation
Table 3 reports the scores obtained by the four LLMs when using all available references for scoring.
SummEval DUC GUMSum
R-1 R-2 R-L BLEU MTR BS BLRT R-1 R-2 R-L BLEU MTR BS BLRT R-1 R-2 R-L BLEU MTR BS BLRT
Qwen 36.43 15.71 31.56 17.41 42.07 89.67 55.15 42.06 17.13 36.67 10.03 31.98 89.96 49.28 44.01 19.23 34.20 22.90 35.25 90.23 47.87
Llama3 37.50 18.09 32.79 22.07 46.09 90.04 56.61 37.44 12.74 32.79 6.45 24.97 89.87 46.01 43.51 19.96 35.34 24.65 36.70 90.33 48.50
Claude 36.03 17.80 31.89 20.84 46.75 89.83 56.07 47.69 20.92 41.73 12.86 39.48 91.27 56.94 44.99 18.65 34.51 20.85 37.62 90.40 50.13
GPT4 35.86 17.90 31.80 21.84 46.36 90.10 57.22 45.73 19.22 39.27 12.07 38.12 91.15 56.01 47.02 19.12 34.23 21.33 43.46 90.17 51.61
Table 3: LLM scores on the multi-reference datasets. R, MTR, BS, and BLRT are short for ROUGE, METEOR, BERTScore, and BLEURT respectively. All available references are used in the evaluation. For ROUGE and BLEURT, we consider the max-variation of the score.
Appendix B Reference-based Metrics
ROUGE.
We use the sacrerouge [^7] https://github.com/danieldeutsch/sacrerouge python implementation of ROUGE Deutsch and Roth (2020), with default parameters.
rouge = Rouge(
max_ngram = 4,
use_porter_stemmer = True,
remove_stopwords = False,
max_bytes = None,
max_words = None,
compute_rouge_l = True,
skip_bigram_gap_length = None,
scoring_function = "max", # or "average"
Notice that the implementation of ROUGE uses the Jackknife method when multiple references are provided.
BLEU.
We use the sacrebleu [^8] https://github.com/mjpost/sacrebleu python implementation of BLEU Post (2018), with default parameters.
bleu = BLEU(
lowercase=False,
force=False,
tokenize=tokenize,
smooth_method=’exp’,
smooth_value=None,
effective_order=False) # True when used at the sentence level
METEOR.
We use the Hugging Face version of Meteor, implemented through the evaluate [^9] https://github.com/huggingface/evaluate library, with default parameters, which wraps the NLTK implementation of the metric. [^10] https://www.nltk.org/api/nltk.translate.meteor_score.html
nltk.translate.meteor_score.meteor_score(references: ~typing.Iterable[~typing.Iterable[str]], hypothesis: ~typing.Iterable[str], preprocess: ~typing.Callable[[str], str] = <method ’lower’ of ’str’ objects>, stemmer: ~nltk.stem.api.StemmerI = <PorterStemmer>, wordnet: ~nltk.corpus.reader.wordnet.WordNetCorpusReader = <WordNetCorpusReader in ’/Users/stevenbird/nltk_data/corpora/wordnet’>, alpha: float = 0.9, beta: float = 3.0, gamma: float = 0.5) -> float[source]
BERTScore.
We use the Hugging Face version of BERTScore, implemented through the evaluate [^11] https://github.com/huggingface/evaluate library, with default parameters, which wraps the bert_score implementation. [^12] https://github.com/Tiiiger/bert_score No TF-IDF weighting is used. Embeddings are obtained by using FacebookAI/roberta-large. We did not fine-tune the model. The corresponding hash is roberta-large_L17_no-idf_version=0.3.12
(hug_trans=4.48.3)
BLEURT.
We use the original Google version of BLEURT. [^13] https://github.com/google-research/bleurt We used the recommended checkpoint,141414available at https://storage.googleapis.com/bleurt-oss-21/BLEURT-20.zip which we did not fine-tune.
Appendix C Variability in Humans and LLMs
Figure 7 compares the variability observed in human-written summaries and in LLM-generated ones. To characterize LLM-generated summaries, given and instance $i$ and a pair of human-written summaries for instance $i$ $R_{ji}$ and $R_{zi}$, we plot $P(o_{i},o_{j})$ where $P$ is the lexical, syntactic, or semantic similarity. To characterize LLM-generated summaries, given and instance $i$ system $S_{j}$ and $S_{z}$, producing outputs $o_{j}$ and $o_{z}$, we plot $P(o_{ji},o_{zi})$.
When compared to summaries generated by different humans, those produced by various LLMs exhibit far less variation, particularly at the semantic and syntactic levels. Quantifying and mitigating the reduced richness and variability of LLM-generated content is an area of ongoing research Guo et al. (2024); Shur-Ofry et al. (2024); Giulianelli et al. (2023) and, while not reflecting poor output quality, raises an open question about whether LLMs can fully replace human references.
(a) human-written summaries
(b) LLM-generated summaries
Figure 7: Variation in human-produced summaries (top) and LLM-generated summaries (bottom).
Appendix D Quality of Scraped References
Figure 8: Models and CNN/DM reference compared to $10$ crowd-sourced references in SummEval. SummEval models are in lighter blue, LLMs in darker blue. The red line indicates the CNN/DM reference score, and the black dotted line represents a random baseline using a different news article as the hypothesis.
One peculiarity of summarization datasets is that they are typically gathered from existing sources Dahan and Stanovsky (2025). This is the case of the CNN/DM dataset Hermann et al. (2015); Nallapati et al. (2016), where articles are scraped from news websites, with the concatenated highlights acting as the target summary. The quality of such targets has been criticized in previous studies Kryscinski et al. (2019); Srivastava et al. (2023), as they contain extraneous facts, references to other articles, and other issues.
To better investigate the impact of these issues—and thus, quantify the gap between scraped summaries and crowd sourced ones—we leverage the additional references in SummEval. To this end, we treat the scraped CNN/DM references as hypotheses, and we score them against the 10 crowd-sourced ones in the SummEval dataset. We assume the latter provide a superior reference set for two main reasons: a) crowd-sourced references are specifically collected to act as summaries, with clear guidelines and an emphasis on quality; b) the larger cardinality of the crowd sourced human references (cf. Table 1) allows for a more comprehensive view of human-produced summaries. We compare these score to those of system outputs.
Figure 8 reports the scores of the original CNN/DM references (red line) and those of the outputs of the summarization systems (blue bars). We also report a random baseline (dotted line) in which a summary of a different document randomly sampled from the collection is used as hypothesis.
Notably, the original CNN/DM references perform worse than all model outputs in all cases but one when evaluated using n-gram-matching metrics. When using BERTScore, the CNN/DM receives a score close to that of the outputs from LLMs. BLEURT rates the original reference higher than SummEval system outputs (except for GPT-2), but still lower than all LLM-generated outputs. These observations corroborate previous concerns on reference quality for summarization, especially when used to score high-quality outputs Goyal et al. (2023).
Appendix E Instance-level variation
Figures 9 and 10 contain the histogram of the instance-level variability for GUMSum and DUC respectively.
Figure 9: Ranges of variability at the instance level on GUMSum. For each instance, we compute the range of the scores of the references scored against the remaining ones.
Figure 10: Ranges of variability at the instance level on DUC. For each instance, we compute the range of the scores of the references scored against the remaining ones.
Appendix F Model Ranking
Table 4 contains the analysis of the rank stability when using one single reference ($k=1$). For SummEval, we considering the cases of ranking the models studied by Fabbri et al. (2021), the LLMs and the combination of the two separately.
Figure 11 shows the Kendall tau correlation between ranks as the number $k$ of references increases. We use the mean version of ROUGE.
R1mean R2mean R3mean R4mean RLmean R1max R2max R3max R4max RLmax BLEU MTR BS BLRT
SummEvalall models min -.04 -.07 .05 -.09 .12 -.06 -.09 -.04 -.06 .04 -.27 .41 .54 .67
avg .51 .56 .57 .51 .63 .51 .56 .56 .52 .62 .38 .76 .79 .86
std .16 .14 .12 .14 .12 .18 .15 .13 .13 .14 .16 .07 .06 .04
max .91 .87 .91 .82 .92 .87 .86 .88 .85 .92 .82 .94 .95 .98
SummEvalpre-LLMs min -.05 -.03 -.05 -.15 0 -.05 -.03 -.12 -.28 .-07 -.30 .57 .28 .50
avg .48 .49 .48 .39 .55 .49 .50 .57 .39 .55 .39 .78 .68 .79
std .14 .13 .13 .15 .13 .13 .13 .15 .16 .12 .17 .06 .09 .07
max .87 .88 .90 .83 .90 .88 .85 .87 .88 .90 .85 .97 .93 .97
SummEvalLLMs min -.33 0 -.33 -1 0 -.33 -.33 .67 -1 0 -.67 0 -67 0
avg .59 .54 .49 .48 .59 .58 .52 .48 .49 .57 .59 .60 .62 .80
std .31 .32 .33 .35 .31 .32 .32 .34 .36 .31 .32 .31 .28 .21
max 1 1 1 1 1 1 1 1 1 1 1 1 1 1
GUMSum min 0 -1 -1 -1 -1 0 -1 -1 -1 -1 -1 0 -1 .33
avg .53 .42 .22 .15 .33 .51 .35 .14 .16 .35 .28 .56 .17 .86
std .32 .40 .44 .47 .41 .32 .43 .47 .47 .43 .45 .31 .48 .19
max 1 1 1 1 1 1 1 1 1 1 1 1 1 1
DUC min .67 .67 .33 -.33 .67 .67 .67 .33 -.33 .67 .33 .67 .33 .67
avg .99 .99 .94 .82 .98 .99 .97 .91 .78 .99 .88 .89 .71 .88
std .05 .05 .12 .22 .08 .05 0.9 .15 .25 .05 .17 .16 .24 .16
max 1 1 1 1 1 1 1 1 1 1 1 1 1 1
Table 4: Rank stability with a single reference. We ranked systems $100$ times and compute the Kendall tau correlation among such rankings. R, MTR, BS, and BLRT are short for ROUGE, METEOR, BERTScore, and BLEURT respectively.
Figure 11: Rank stability when increasing the number of references over all three datasets. For ROUGE, we show the mean variant. Notice we use different ranges for the y axes for each dataset to improve readability.
Appendix G System-level Correlation
Figure 12 shows the instance-level correlation for SummEval (top) and GUMSum (bottom) using ROUGEavg.
Figure 13 show the system-level correlation on SummEval with ROUGEmax (top) and ROUGEavg (bottom).
Figure 12: Pearson correlation at the instance level on SummEval (top) and GUMSum (bottom) using ROUGEmean.
Figure 13: Kendall tau at the system level on SummEval using ROUGEmax (top) and ROUGEmean (bottom).